OpenStudy (anonymous):

5!·4! is equivalent to:

OpenStudy (anonymous):

5! = 5x4x3x2x1 what 4!??

OpenStudy (anonymous):

\[5!\times4!=(1\times 2\times 3\times 4\times 5)\times(1\times 2\times 3\times 4)\]

OpenStudy (anonymous):

so then whats the answer

OpenStudy (anonymous):

(5!)(4!)² (5)(4!)² 20! these are my options is it 20!

OpenStudy (skullpatrol):

No.

OpenStudy (anonymous):

Use a calculator.

OpenStudy (anonymous):

i did how would you do this ! on a calculator

OpenStudy (anonymous):

Test all the options with your nifty calculator.

OpenStudy (anonymous):

Your calculator should have it. You should really get an approved calculator.

OpenStudy (anonymous):

In Australia we're required to have the "Board of Studies" requirement of a calculator.

OpenStudy (agent0smith):

Pull common factors from Azteck's answer: \[5!\times4!=(1\times 2\times 3\times 4\times 5)\times(1\times 2\times 3\times 4) = 5 \times (1 \times 2 \times 3 \times 4)^{2}\]

jimthompson5910 (jim_thompson5910):

Hint: 5! = 5*4!

OpenStudy (agent0smith):

^ that way makes it easier.

jimthompson5910 (jim_thompson5910):

so that means 5! * 4! is the same as 5*4! * 4!

OpenStudy (agent0smith):

There's no need for a calculator here, it probably makes it more work.

OpenStudy (anonymous):

i have one of those 200 dollar calculators it doesnt have it

jimthompson5910 (jim_thompson5910):

if you have a TI calculator, then you can use the ! or factorial button to get the job done

OpenStudy (anonymous):

2880

OpenStudy (skullpatrol):

Let her work it out by hand first.

jimthompson5910 (jim_thompson5910):

if you have a TI 83 for instance, you hit the math button, then you go to the probability menu, then scroll down to 4

OpenStudy (anonymous):

i got 2880 but thats not an option

OpenStudy (skullpatrol):



OpenStudy (anonymous):